PART V: Recall that for scalar > 0, the probability density function of an exponential random variable with parameter , isfor which 12) (1 point] What is the maximum likelihood estimator? In other words, what is the value of the derivative of (D;)
Answer #1

_P (x, d) =_Ae-dx. for realisations D={4,2 ... ? ___CID, A). - P (24), २). Hि ___- 2 - 4 hence log LCD id) __ = one (R,A)

Note that maximum likelihood estimation is nothing but the maximisation of the likelihood function.

hence loge = nloga -. a I x, 2 loge = 0 Sample meno - Hente 1 2 1 2 logh = - 1 / 2 <0. loge is maximised